OpenStudy (falling_in_katt):

The geometric mean of 8 and 2 is 4. True or False

OpenStudy (mathmate):

hint: geometric mean of x1 and x2 is equal to \(\sqrt{x1\times x2}\) So you can verify what the geometric mean of 2 and 8 should be.

OpenStudy (mathstudent55):

x is the geometric mean of a and b if a is to x as x is to b, or \(\dfrac{a}{x} = \dfrac{x}{b} \) Then when you take the cross products and set them equal, you get: \(x^2 = ab\) Solving for x, you get: \(x = \sqrt{ab} \)

OpenStudy (mathstudent55):

In your problem, use 8 and 2 as a and b, and solve for x.

OpenStudy (falling_in_katt):

True?

OpenStudy (mathstudent55):

yes